OpenStudy (♥ashley♥):

Mr. Jefferson has 20 students in his class. He wants to send 4 of these students to pick up books for the class. How many combinations of 4 students can he choose? A. 5 B. 80 C. 4845 D. 116,280

OpenStudy (anonymous):

this uses a formula for allocations where the order does not matter \[which is A = \frac{ n! }{(n-m)! }\] so 20! / 16! = 17 * 18 * 19 * 20 = 116280

OpenStudy (♥ashley♥):

Thanks!

OpenStudy (math&ing001):

I think the number of combination is calculated like this : \[C _{20}^{4}=\frac{ n! }{ (n-r)!*r! }=\frac{ 20! }{ 16!*4! }=4845\]